OpenStudy (anonymous):

Ken has a pole that is 26 inches long. He wishes to cut it into two pieces so that one piece will be 8 inches longer than the other. How long must the shorter piece be?

OpenStudy (beccaboo333):

\[26\div2\] Then subtract 8 from one of those two, you have your answer.

OpenStudy (warriorz13):

I am guess 5 because 28/2=13-8=5 Hope it helps

OpenStudy (beccaboo333):

Please don't give the answer ._. Just the problem on how to get the answer.. Always need our users working as much as us.. @Warriorz13

OpenStudy (anonymous):

Lets call the total length 't', the short piece 's' and the long piece 'l'. We know the following: \[t=s+l\] and \[l=s+8\] so by substitution:\[26=s+s+8=2s+8\] So \[s=9\] and \[l=s+8=9+8=17\] The long piece is 17cm the short piece is 9cm for a total of 26cm. We can check our answer: 17-9=8cm longer, just like we wanted.

OpenStudy (warriorz13):

OK NP Will try to remember that next time